OpenStudy (anonymous):

how would you graph in a circle (x-2)^2 + (y+1)^2 =9?

OpenStudy (superdavesuper):

hi, katie - similar Q to the other one u just worked on but this time u got the formula. so what is the radius?

OpenStudy (anonymous):

is 9 not the radius?

OpenStudy (noseboy908):

No, the formula your original equation is given in is "(x-x1)+(y-y1)=r^2". So, seeing that your given number for "r^2" is nine, what is r?

OpenStudy (noseboy908):

Oh, and the parenthesis should be squared, but everything else is correct in it.

OpenStudy (anonymous):

1^2 +1^2? idk how to solve im confused on where to start?

OpenStudy (noseboy908):

Try not to over complicate it, the "squared" that goes outside the sets of parenthesis is only for form, to tell that it's a circle. It isn't actually used to solve anything right now. The only important parts of the equation is what is being added or subtracted from "x" and "y", and what it all is equal to. So, I'd suggest starting by finding r. R is found here by simply square-rooting the number everything is equal to. So, what is r?

OpenStudy (anonymous):

9^2 = 81

OpenStudy (noseboy908):

No, remember to square root the number, don't square it. Try again.

OpenStudy (anonymous):

so square root 9? 3?

OpenStudy (noseboy908):

Yep! Remember that: When given the form "(x-x1)^2 + (y=y1)^2 = r^2", you always have to SQUARE ROOT the "r^2" in order to get your radius. Now, knowing that 3 is your radius, let's find your center. Please restate the equation as given left of the equal sign.

OpenStudy (anonymous):

which equation?